# Non-local, diamagnetic electromagnetic effects in long MITLs

### v1

@@ Abstract
We identify the physics responsible for the critical reduction of current losses in magnetically insulated transmission lines (MITLs) before magnetic (self-) insulation has been established. Focusing on time-dependent physics, a drive current with a prototypical sine squared temporal profile introduces sufficiently strong time dependence that steady state results alone become insufficient for the complete understanding of current losses. We find that the effects of time dependence are most pronounced in long MITLs, i.e., MITLs in which an electromagnetic wave traverses its length in time comparable to the current pulse length. The time-dependent physics can be described as non-local, diamagnetic electromagnetic response of space charge limited currents. As the length of the MITL is reduced (in the above sense), current losses converge to those based on the well known Child-Langmuir law in static external fields. We present a simple one-dimensional (1D) model that encapsulates the essence of this physics. We find excellent agreement with 2D particle-in-cell (PIC) simulations for two MITL geometries, Cartesian parallel plate and azimuthally symmetric straight coaxial. Based on the 1D model, we explore various scaling dependencies of current losses with relevant parameters, e.g., peak current, peak pulse time, geometrical dimensions, etc. We also propose an improved physics model of magnetic insulation, which could help improve predictions of current losses by common circuit element codes, such as BERTHA. Lastly, we describe how to calculate temperature rise due to electron impact, as a diagnostic within the 1D model.
magnetically insulated transmission line; MITL; parallel plate; coaxial; magnetic insulation; non-local; diamagnetic; particle-in-cell; PIC; kinetic; space charge limited; SCL; electron emission; BERTHA; circuit element model; $Z$ machine.
@@ I Introduction
Pulsed power devices sinars_review_2020 ; McBride:2018 ; stygar_design_2015 require the transport of large amounts of power over relatively long distances. Previous experiments baranchikov_transfer_1977 ; di_capua_propagation_1979 ; vandevender_long_1979 have shown that $80$% or more of the input energy can be delivered to a load up to 101010\,10m away from a generator. At Sandia National Laboratories’ $Z$ machine, up to 404040\,40terawatts of power is delivered from the stack to a load over more than a meter distance, with little losses. At the heart of the ability to transport such large amounts of power is the physics of magnetic insulation (also called self-insulation) in magnetically insulated transmission lines (MITLs) lovelace_theory_1974 ; Creedon:1975 ; di_capua_magnetic_1983 ; Mendel:1983 ; Ottinger:2006 .
During the beginning of the current pulse, electron emission from the cathode is observed when the electric fields exceed values of about 20202020303030\,30MV/m di_capua_propagation_1979 . These electrons may or may not be initially insulated, something highly dependent on MITL geometry and temporal current pulse profile. For the prototypical sine squared current pulse shape used throughout this work,  [^1] The conclusions in this work also hold for more realistic current pulses used on $Z$, however, we do not show results for such pulses. uninsulated electrons are typically observed in long MITLs, i.e., MITLs in which an electromagnetic (EM) wave traverses its length in time comparable to the length of the current pulse itself. The longer the MITL, the longer the time interval of uninsulated emission. Uninsulated electrons may contribute significantly to current loss and anode temperature rise and are for this reason a major concern when designing long MITLs.
Current loses are being predicted in one of two ways. First, fully kinetic electromagnetic particle-in-cell (PIC) simulations are considered as the most general and trusted approach pointon_particle_cell_2007 ; pointon_current_2009 ; ottinger_rescaling_2006 ; welch_electrode_2019 ; bennett_current_2019 ; welch_fast_2020 . By its nature, the PIC method birdsall_plasma_2005 ; hockney_computer_2021 is very computationally expensive. For this reason, approximate one-dimensional (1D) circuit element models (CEM) have been devised hinshelwood_bertha_1983 ; huttlin_tlines_1983 ; weseloh_tlcode_1989 ; harjes_circuit_2003 ; spielman_design_2019 . Since CEM models are only 1D, their computational cost is many orders of magnitude lower than that of PIC, and for this reason are preferred for preliminary MITL design, leaving PIC simulations for final verification. Both methods have shown reliable predictions of current losses in time-dependent settings (i.e., using current pulses), implying that they both capture the essential physics.
@@
    SCL currents vary according to the well known Child-Langmuir law child_discharge_1911 ; langmuir_effect_1913 ;
@@
    The magnetic insulation ramps down SCL currents smoothly, according to a Hull curve. [^2] We use this term in the hope that there is no confusion with a similar one used in the published literature.
In the main exposition, we consider a parallel plate Cartesian geometry, which allows for a certain degree of analytical treatment. The straight coaxial geometry with a finite AK gap can also be treated analytically; however, its treatment greatly simplifies in the small AK gap limit, for which we outline some detail in Appendix A.
The paper is structured as follows. Section I is an introduction. Section II introduces the reduced 1D electromagnetic model. Section III describes the non-local, diamagnetic time-dependent effects in long MITLs. Section IV presents scaling of MITL losses with various parameters. Section V presents a model for magnetic insulation, improving upon a similar such in CEM models. Section VI presents the temperature rise diagnostics within the 1D model. Section VII summarizes and presents final concluding remarks.
@@ II A one-dimensional electromagnetic model of a MITL
In the geometry under consideration, all electromagnetic waves, external and emitted by SCL currents, propagate as TEM modes. We choose $z$ as the propagation direction, $x$ as the transverse (AK gap) direction, then $y$ is the ignorable direction (symmetry dimension). Correspondingly, the non-trivial components of the fields and currents (current densities) are $E_{x}$, $B_{y}$, and $j_{x}$. Quantities also depend on time, $t$. Waves are launched at $z=0$ in the positive $z$-direction. Waves propagating in the negative $z$-direction (to the left) are absorbed at $z=0$. There is a perfectly conducting load (short) at the opposite end of the MITL, $z=L$, so that waves are reflected off that end.
The basic model of SCL emission assumes that a local SCL current is “once-on-always-on.” In other words, we assume that once the threshold for SCL emission is exceeded at a particular location and instant of time, the inventory of plasma is sufficient to supply any amount of electrons necessary to sustain zero electric field at the cathode at maximal SCL current; this is the typical assumption in PIC as well. Some models assume initially a source-limited emission, which later transitions into SCL emission. We do not include such detail as it does not qualitatively affect the following discussion and results.
As already mentioned, the small AK gap assumption allows us to decouple the two spatial dimensions of the problem. The two 1D problems consist of an electromagnetic problem along the direction of EM wave propagation and an electrostatic problem in the direction across the AK gap. First, let us consider the electromagnetic part of the model. Suppose a spatial location $z_{0}$ starts emitting SCL current at time $t_{0}=0$. This emission model can be mathematically represented as a Dirac delta-function in space and a Heaviside step-function in time
@@
with current $I_{0}$ per unit length (unit length in the ignorable $y$-direction). We assume $I_{0}$ to be the current on the conductor. The relation between the current in the AK gap and the current on the conductor is illustrated in Fig. 1. We also note that the delta function has dimensions of inverse length while the step function is dimensionless; thus $j_{x}$ has the dimensions of current density, e.g., Ampere/m2 in SI units.
Working in the Coulomb gauge, the equation for the electromagnetic waves can be reduced to the wave equation for the vector potential
@@
Using the fundamental solution (Green’s function) of the wave operator (d’Alembertian), $\partial_{t}^{2}-c^{2}\partial_{x}^{2}$, in one spatial dimension (of infinite $z$-extent), ${\cal E}_{1}=\frac{1}{2c}\theta\left(t-\frac{|z|}{c}\right)$, we follow the general rule for finding the special solution of Eq. (2) by convolution vladimirov_equations_1984 , $A_{x}(z,t)={\cal E}_{1}\star\left(c^{2}\mu_{0}j_{x}\right)$. The integrals are easy to evaluate and give
@@
where $\epsilon_{0}$ and $\mu_{0}$ are the permittivity and permeability of vacuum, $Z_{0}=\sqrt{\mu_{0}/\epsilon_{0}}\approx 376.7\,\Omega$ is the vacuum (free space) impedance, and $c$ is the speed of light in vacuum. The fields following from Eqs. (3), (4) are
@@
$B_{y}(z,t)$ $=-\operatorname{sign}(z-z_{0})\,\frac{E_{x}(z,t)}{c}.$ (7)
@@
Note that the magnetic field has a discontinuity at the current location, $z=z_{0}$, since this is a surface (sheet) current. The jump across the discontinuity is the difference between the values on the right and on the left of the current sheet, and gives the expected result
$\left[B_{y}\right]_{z=z_{0}}=B_{y}(0^{+})-B_{y}(0^{-})=\frac{I_{0}Z_{0}}{c}=\mu_{0}I_{0}.$ (8)
@@
An external (vacuum) field can be launched into the MITL by setting up a current source at the launch side of the MITL, say $z_{0}=0$, with a desired temporal dependence. For example, a sine squared current pulse peaking at value $I_{\rm peak}$ at time $t=\tau_{\rm peak}$ can be set up by
$j_{x}(z,t)=-I_{\rm peak}\,\delta(z)\,\sin^{2}\left(\frac{\pi t}{2\tau_{\rm peak}}\right).$ (9)
The corresponding fields can be calculated by following the steps leading to Eq. (5):
$E^{\rm ext}_{x}(z,t)$ $=\frac{I_{\rm peak}Z_{0}}{2}\,\sin^{2}\left(\frac{\pi(t-|z|/c)}{2\tau_{\rm peak}}\right),$ (10)
$B^{\rm ext}_{y}(z,t)$ $=\frac{E_{x}(z,t)}{c},$ (11)
@@
where the magnetic field sign corresponds to a wave launched in the positive $z$-direction. Our numerical implementation indeed uses Eq. (9) to launch external waves into the MITL.
@@
Now consider the dynamics within the AK gap. At any instant of time, we assume the SCL current is given by Child-Langmuir’s law,
$j_{x}(z,t)\equiv j_{\rm CL}=\frac{4\epsilon_{0}}{9d^{2}}\sqrt{\frac{2e}{m}}V(z,t)^{3/2},$ (12)
where $d$ is the size of the AK gap, $e$ and $m$ are the charge and mass of the electron, and $V(z,t)$ is calculated from the electric field Eq. (6) as follows. Within the AK gap, we know from the steady state Child-Langmuir law (see for example davidson_physics_2001 ) that at any given $z$ and $t$
$V(x;z,t)=V_{0}\left(\frac{x}{d}\right)^{4/3},$ (13)
where $x=0$ is the cathode and $x=d$ is the anode. The semicolon in the list of arguments in Eq. (13) emphasizes the fact that $z$ and $t$ are here considered as parameters, not actual arguments to $V$. At the anode, the total electric field is the sum of the external (electromagnetic) field (6) and the space charge field:
$E_{x}^{\rm tot}(x=d)=-\left.\frac{dV(x)}{dx}\right|_{x=d}=-\frac{4}{3}\frac{V_{0}}{d}.$
We now identify the value of the electric potential at the anode at any $z$ and $t$ with the voltage at the anode, $V_{0}=V(z,t)=-E_{x}(z,t)\,d$. This gives the potential determining the value of the current $j_{x}(z,t)$ in Eq. (2),
$j_{x}(z,t)=\frac{4\epsilon_{0}}{9d^{2}}\sqrt{\frac{2e}{m}}\left[E_{x}(z,t)\,d\right]^{3/2},$ (14)
completing the system of equations (2)–(4). (Absolute values are taken whenever necessary.)
@@
The total electric field at the anode is given by
$E_{x}^{\rm tot}(x=d;z,t)=\frac{4}{3}E_{x}(z,t),$ (15)
@@
$V_{0}=V(z,t)=\frac{3}{4}E_{x}(z,t)\,d.$ (16)
We use equations (15) and (16) in two cases. First, in the 1D code output, we augment the electric field found from Eq. (6) by $4/3$, according to Eq. (15), at any location where uninsulated SCL emission has initiated. For definiteness, we compare fields and currents between models near the anode. And second, relation (16) is instrumental to correctly determine the Hull curve from PIC simulations, as discussed in Sec. V.
As a final comment in this section, we note that the simulation of SCL emission by PIC has both physical and numerical aspects. The physical aspect is that of the Child-Langmuir law. In emission due to field stress, after a certain threshold value of the (normal) electric field has been exceeded, electrons are emitted from the cathode and conduct space charge limited current in the anode-cathode gap, when uninsulated. The numerical aspect aims to implement an algorithm that uses computational particles and finite time steps, that is numerically stable, and that produces results in agreement with the physical aspect. In this regard, “numerical knobs” exist in PIC codes which, if not used carefully, may noticeably distort the results. For example, Sandia’s PIC code EMPIRE bettencourt_empire-pic_2021 , implements a function that ramps up SCL current in time from a small fraction to its full value according to some functional dependence. A similar capability exists in CHICAGO welch_implementation_2004 ; welch_adaptive_2007 ; genoni_fast_2010 ; welch_fast_2020 . We believe that most PIC codes allow for similar numerical knobs. Accordingly, we have implemented a similar parameter our 1D electromagnetic model, allowing for a closer comparison of the results. We refer the reader to the extensive literature for more detail (e.g., birdsall_plasma_2005 ; pointon_slanted_1991 ; watrous_improved_2001 ; luginsland_beyond_2002 , with a brief discussion also given in Ref. evstatiev_efficient_2023 ).

### v2

# Non-local, diamagnetic electromagnetic effects in magnetically insulated transmission lines

@@ Abstract
We identify [^1] Copyright (2025) E. G. Evstatiev, M. H. Hess, N. D. Hamlin, and B. T. Hutsel. This article is distributed under a Creative Commons Attribution-NonCommercial-NoDerivs 4.0 International (CC BY-NC-ND) License. This article appeared in Phys. Plasmas 32, 062707 (2025) and may be found at https://doi.org/10.1063/5.0254084. the time-dependent physics responsible for the critical reduction of current losses in magnetically insulated transmission lines (MITLs) due to uninsulated space charge limited (SCL) currents of electrons emitted by field stress. A drive current of sufficiently short pulse length introduces a strong enough time dependence that steady state results alone become inadequate for the complete understanding of current losses. The time-dependent physics can be described as a non-local, diamagnetic electromagnetic response of space charge limited currents. As the pulse length is increased or equivalently, the MITL length reduced, these time-dependent effects diminish and current losses converge to those predicted by the well-known Child-Langmuir law in the external (vacuum) fields. We present a simple one-dimensional (1D) model that encapsulates the essence of this physics. We find excellent agreement with 2D particle-in-cell (PIC) simulations for two MITL geometries, Cartesian parallel plate and azimuthally symmetric straight coaxial. Based on the 1D model, we explore various scaling dependencies of MITL losses with relevant parameters, e.g., peak current, pulse length, geometrical dimensions, etc. We propose an improved physics model of magnetic insulation in the form of a Hull curve, which could also help improve predictions of current losses by common circuit element codes, such as BERTHA. Lastly, we describe how to calculate temperature rise due to electron impact within the 1D model.
electron emission; space charge limited; SCL; Child-Langmuir; time-dependent; electromagnetic; non-local; diamagnetic; magnetic insulation; Hull curve; magnetically insulated transmission line; MITL; parallel plate; coaxial; particle-in-cell; PIC; kinetic; 1D; 2D; BERTHA; circuit element; $Z$ machine; pulsed power; power flow.
@@ I Introduction
Pulsed power devices [1, 2, 3] require the transport of large amounts of power over relatively long distances. Previous experiments [4, 5, 6] have shown that $80$% or more of the input energy can be delivered to a load up to 101010\,10m away from a generator. At Sandia National Laboratories’ $Z$ machine, up to 404040\,40terawatts of power is delivered from the stack to a load over more than a meter distance, with little losses. At the heart of the ability to transport such large amounts of power is the physics of magnetic insulation (also called self-insulation) in magnetically insulated transmission lines (MITLs) [7, 8, 9, 10, 11, 12].
During the beginning of the current pulse, electron emission from the cathode is observed when the electric fields exceed values of about 20202020303030\,30MV/m [5]. These electrons may or may not be initially insulated, something highly dependent on MITL geometry and temporal current pulse profile. For the prototypical sine squared current pulse shape used throughout this work,  [^2] The conclusions in this work also hold for more realistic current pulses used on $Z$, however, we do not show results for such pulses. uninsulated electrons are typically observed in long MITLs (see the definition below). Uninsulated electrons are a major contributor to current loss and anode temperature rise, and are for this reason a major concern when designing long MITLs. Additionally, when the anode temperature rises above approximately 400 ${}^{\circ}\!\!$ C, contaminants are thermally desorbed from the electrode surface [13], becoming a source of ions at the anode. Ion Larmor radii are typically larger than those of electrons and for this reason, in many situations ions can account for a significant fraction of the current loss. Ion current losses are not treated in this work but will be considered in the future.
Current losses are being predicted in one of two ways. First, fully kinetic electromagnetic (EM) particle-in-cell (PIC) simulations are considered as the most general and trusted approach [14, 15, 16, 17, 18, 19]. By its nature, the PIC method [20, 21] is very computationally expensive. For this reason, approximate one-dimensional (1D) circuit element models (CEM) have been devised [22, 23, 24, 25, 26, 27]. Since CEM models are only 1D, their computational cost is many orders of magnitude lower than that of PIC, and for this reason are preferred for preliminary MITL design, leaving PIC simulations for final verification. Both methods have shown reliable predictions of current losses in time-dependent settings (i.e., using current pulses), implying that they both capture the essential physics.
@@
    SCL currents vary according to the well known non-relativistic Child-Langmuir law [28, 29];
@@
    A magnetic insulation model, in which SCL currents ramp down smoothly as the magnetic field increases, according to what we refer to as a Hull curve (see Sec. V).
The “standard” definition of a long MITL is one in which an electromagnetic wave traverses its length in time comparable to the length of the electromagnetic (current) pulse itself. However, based on the results in this work, we propose an alternative definition of a long MITL, which more accurately accounts for the electromagnetic effects and more generally, the effects of time dependence.
Definition.
A long MITL is one in which time-dependent processes result in a significant deviation of quantities of merit from their steady state values.
Relevant quantities of merit include fields, current and charge densities, loss current and charge, anode temperature rise, etc. An example considered in this work is a 100100100\,100ns pulse launched into a 111\,1m-long MITL. Although the travel time (back and forth) of the EM wave in this MITL is only 6.676.676.67\,6.67ns, much smaller than the pulse length of 100100100\,100ns, we find that current losses are greatly overestimated by using steady-state theory alone. This and many similar results have served as the motivation for the above definition.
In the main exposition, we consider a parallel plate Cartesian geometry, which allows for a certain degree of analytical treatment. The straight coaxial geometry with a finite AK gap can also be treated analytically; however, its treatment greatly simplifies in the small AK gap limit, for which we outline some detail in Appendix A. In general geometries, we expect the physical nature of the processes to be the same, albeit with quantitative differences.
The paper is structured as follows. Section I is an introduction. Section II introduces the reduced physics 1D electromagnetic model. Section III describes the non-local, diamagnetic time-dependent electromagnetic effects in long MITLs. Section IV presents scaling of MITL losses with various parameters. Section V presents a model for magnetic insulation, improving upon those found in present CEM models. Section VI presents the temperature rise diagnostics within the 1D model. Section VII summarizes and presents final concluding remarks.
@@ II A one-dimensional electromagnetic model of a MITL
In the geometry under consideration, all electromagnetic waves, external and emitted by SCL currents, propagate as TEM modes. We choose $z$ as the propagation direction, $x$ as the transverse (AK gap) direction, then $y$ is the ignorable direction (symmetry dimension). Correspondingly, the non-trivial components of the fields and currents (current densities) are $E_{x}$, $B_{y}$, $j_{x}$. Quantities also depend on time, $t$. Waves are launched at $z=0$ in the positive $z$-direction. Waves propagating in the negative $z$-direction (to the left) are absorbed at $z=0$. There is a perfectly conducting load (short) at the opposite end of the MITL, $z=L$, so that waves are reflected off that end.
The basic model of SCL emission assumes that a local SCL current is “once-on-always-on.” In other words, we assume that once the threshold for SCL emission is exceeded at a particular location and instant of time, the inventory of plasma is sufficient to supply any amount of electrons necessary to sustain zero electric field at the cathode at maximal SCL current; this is the typical assumption in PIC simulations as well. Some models assume initially a source-limited emission, which later transitions into SCL emission. We do not include such detail as it does not qualitatively affect the following discussion and results.
As already mentioned, the small AK gap assumption (A-i) allows us to decouple the two spatial dimensions of the problem. The two 1D problems consist of an electromagnetic problem along the direction of EM wave propagation and an electrostatic problem in the direction across the AK gap. First, let us consider the electromagnetic part of the model. Suppose a spatial location $z_{0}$ starts emitting an SCL current at time $t_{0}=0$. This emission model can be mathematically represented as a Dirac delta-function in space and a Heaviside step-function in time
@@
with current $I_{0}$ per unit length (unit length in the ignorable $y$-direction). We assume $I_{0}$ to be the current on the conductor. The relation between the current in the AK gap and the current on the conductor is illustrated in Fig. 1. We also note that the delta function has dimensions of inverse length while the step function is dimensionless; thus $j_{x}$ has the dimensions of current density, e.g., A/m2 in SI units.
Working in the Coulomb gauge (vector potential satisfying $\nabla\cdot\mathbf{A}=0$), the equation for the electromagnetic waves can be reduced to
@@
Using the fundamental solution (Green’s function) of the wave operator (d’Alembertian), $\partial_{t}^{2}-c^{2}\partial_{x}^{2}$, in one spatial dimension (of infinite $z$-extent), ${\cal E}_{1}=\frac{1}{2c}\theta\left(t-\frac{|z|}{c}\right)$, we can follow the general rule for finding the special solution of Eq. (2) by convolution [30], $A_{x}(z,t)={\cal E}_{1}\star\left(c^{2}\mu_{0}j_{x}\right)$. For the current source (1), the integrals are easy to evaluate and give
@@
where $\epsilon_{0}$ and $\mu_{0}$ are the permittivity and permeability of vacuum, $Z_{0}=\sqrt{\mu_{0}/\epsilon_{0}}\approx 376.7\,\Omega$ is the vacuum (free space) impedance, and $c$ is the speed of light in vacuum. The fields following from Eqs. (3)–(5) are
@@
$B_{y}(z,t)$ $=\operatorname{sign}(z-z_{0})\,\frac{E_{x}(z,t)}{c}.$ (7)
@@
Note that the magnetic field has a discontinuity at the location, $z=z_{0}$, since this is a surface (sheet) current. The jump across the discontinuity is the difference between the values on the right and on the left of the current sheet, and gives the expected result
$\left[B_{y}\right]_{z=z_{0}}=\left|B_{y}(0^{+})-B_{y}(0^{-})\right|=\frac{I_{0}Z_{0}}{c}=\mu_{0}I_{0}.$ (8)
@@
An external (vacuum) field can be launched into the MITL by setting up a current source at the launch side of the MITL, $z_{0}=0$, with a desired temporal dependence. For example, a sine squared current pulse peaking at value $I_{\rm peak}$ at time $t=\tau_{\rm peak}$ can be set up by
$j_{x}^{\rm ext}(z,t)=I_{\rm peak}\,\delta(z)\,\sin^{2}\left(\frac{\pi t}{2\tau_{\rm peak}}\right).$ (9)
The corresponding fields can be calculated by following the same steps leading to Eq. (5):
$E^{\rm ext}_{x}(z,t)$ $=-\frac{I_{\rm peak}Z_{0}}{2}\,\sin^{2}\left(\frac{\pi(t-|z|/c)}{2\tau_{\rm peak}}\right),$ (10)
$B^{\rm ext}_{y}(z,t)$ $=\frac{{E^{\rm ext}_{x}(z,t)}}{c},$ (11)
@@
where the magnetic field sign corresponds to a wave launched in the positive $z$-direction (see Fig. 1). Our numerical implementation indeed uses Eq. (9) to launch external waves into the MITL.
@@
Now consider the dynamics within the AK gap. Assuming the (vacuum) electric field at the cathode has exceeded the SCL emission threshold, at any instant of time we assume that the SCL current is given by Child-Langmuir’s law,
${j_{\rm SCL}(z,t)=\frac{4\epsilon_{0}}{9d^{2}}\sqrt{\frac{2e}{m}}V(z,t)^{3/2}},$ (12)
[…]
${j_{\rm SCL}(z,t)=\frac{4\epsilon_{0}}{9d^{2}}\sqrt{\frac{2e}{m}}\left[E_{x}(z,t)\,d\right]^{3/2}.}$ (13)
(Absolute values are taken whenever necessary.) Eq. (13) is seen as Assumption (A-ii). The final expression for the current density is obtained by applying Assumption (A-iii) with the help of an analytic fit of PIC simulation data, as discussed in Sec. V. Thus, the current density entering the right-hand side of Eq. (2) is given by
${j_{x}(z,t)=Y\!\left(\frac{B_{y}(z,t)}{B_{\rm crit}(z,t)}\right)\,j_{\rm SCL}(z,t),}$ (14)
where the function $Y$ is given in Eq. (21), $B_{y}$ is given by (4), and $B_{\rm crit}$ is given by (20). The system of equations (2)–(4), (13), (14) is complete. We have shown explicitly all arguments in Eq. (14) in order to emphasize the fact that all quantities defining our 1D model are local in space and time. It is also assumed that when the electric field at the cathode surface is positive, i.e., has the “wrong” sign, no SCL current is emitted.
In the presence of uninsulated SCL emission at $(z,t)$, the electric field within the AK gap is modified and is no longer constant. The Child-Langmuir condition of zero electric field at the cathode is satisfied by (e.g., see [31])
$V(x;z,t)=V_{0}\left(\frac{x}{d}\right)^{4/3}.$ (15)
The semicolon in the list of arguments in Eq. (15) emphasizes the fact that $z$ and $t$ are here considered as parameters, not actual arguments to $V$. The modified electric field — denote that by $\widetilde{E}_{x}(x;z,t)$ — is found from (15) and at the anode has the value (for definiteness, we compare fields and currents between models near the anode)
$\widetilde{E}_{x}(x=d)=-\left.\frac{dV(x)}{dx}\right|_{x=d}=-\frac{4}{3}\frac{V_{0}}{d}.$
@@
Therefore, the electric field at the anode in the presence of uninsulated SCL emission is given by
$\widetilde{E}_{x}(x=d;z,t)=\frac{4}{3}E_{x}(z,t),$ (16)
@@
${V_{0}=V(z,t)=\frac{3}{4}\widetilde{E}_{x}(z,t)\,d.}$ (17)
We use equations (16) and (17) in two cases. First, Eq. (16) allows us to find the electric field at the anode in the case of uninsulated SCL emission from the solution of the 1D model, $E_{x}(z,t)$, by simply multiplying it by $4/3$; conversely, this factor is not included at any $(z,t)$ prior to initiation of SCL emission or after magnetic insulation has established. The electric field $\widetilde{E}_{x}(z,t)$ can be directly compared to its counterpart from a full electromagnetic PIC simulation. Keep in mind that this is a dynamic system and at the same location $z$, even after SCL emission has once started, magnetic insulation can repeatedly cease and at a later time reestablish. And second, relation (17) is instrumental to correctly determine the Hull curve from PIC simulations, as discussed in Sec. V. For the remainder of the paper, we will denote the electric field in the system by $E_{x}(z,t)$, implicitly including the factor $4/3$ whenever necessary.
Numerical considerations
We note that the simulation of SCL emission by PIC has both physical and numerical aspects. The physical aspect is that of the Child-Langmuir law. In emission due to field stress, i.e., after a certain threshold value of the (normal) electric field has been exceeded, electrons are emitted from the cathode and conduct space charge limited current in the anode-cathode gap, when uninsulated. The numerical aspect aims to implement an algorithm that uses computational particles and finite time steps, that is numerically stable, and that produces results in agreement with the physical aspect. In this regard, “numerical knobs” exist in PIC codes which, if not used carefully, may noticeably distort the results. For example, Sandia’s PIC code EMPIRE [32], implements a function that ramps up SCL current in time from a small fraction to its full value, according to some temporal dependence. A similar capability exists in CHICAGO [33, 34, 35, 19]. We believe that most PIC codes allow for similar numerical knobs. Accordingly, we have implemented a similar parameter in our 1D electromagnetic model, allowing for a closer comparison of the results. Typical ramp times used in our simulations were in the 0.250.250.250.250.50.50.5\,0.5ns range. We refer the reader to the extensive literature for more detail (e.g., [20, 36, 37, 38], with a brief discussion also given in Ref. [39]).
[…]